# Embedding in Recommender Systems: A Survey

### v2

@@ 8. Conclusion
In summary, the use of embeddings in recommender systems has shown great promise in recent years. These methods can capture complex relationships between items and users, leading to better performance compared to traditional approaches. However, there are still challenges to overcome, such as scalability and understandability. This survey aims to provide a comprehensive overview of how embeddings are used in recommender systems, including the different techniques employed, their effectiveness, and the current issues and future directions in the field. By sharing recent advancements, this survey paves the way for future exploration and improvements in the ever-evolving landscape of recommendation systems.

### v3

@@ 9. Conclusion
In summary, the use of embeddings in recommender systems has shown great promise in recent years, demonstrating significant advantages in capturing complex relationships between users and items. From traditional matrix factorization approaches to advanced sequential modeling and graph-based methods, embedding techniques have evolved to address various challenges in recommendation scenarios. The integration of self-supervised learning has further enhanced embedding quality without requiring extensive labeled data, while practical considerations have led to innovations in embedding optimization through AutoML, hashing, and quantization techniques. The recent emergence of large language models has opened new possibilities for enhancing embedding quality through semantic understanding and knowledge transfer, though challenges remain regarding computational efficiency and practical deployment. However, there are still several critical challenges to overcome, including scalability issues, the need for better interpretability, and concerns about fairness and bias in learned representations. This survey provides a comprehensive overview of embedding techniques in recommender systems, covering their theoretical foundations, practical applications, and effectiveness across different scenarios. By systematically examining current methodologies and highlighting future research directions, this work aims to serve as a valuable resource for researchers and practitioners in the field. As recommender systems continue to evolve and tackle more complex scenarios, embedding techniques will remain crucial for their success, and the insights shared in this survey pave the way for future innovations and improvements in this rapidly developing landscape.
Acknowledge
This research was partially supported by National Natural Science Foundation of China (No.62502404), Hong Kong Research Grants Council’s Research Impact Fund (No.R1015-23), Research Grants Council’s Collaborative Research Fund (No.C1043-24GF), Research Grants Council’s General Research Fund (No.11218325), Institute of Digital Medicine of City University of Hong Kong (No.9229503), Huawei (Huawei Innovation Research Program), Tencent (CCF-Tencent Open Fund, Tencent Rhino-Bird Focused Research Program), Alibaba (CCF-Alimama Tech Kangaroo Fund No. 2024002), Ant Group (CCF-Ant Research Fund), Didi (CCF-Didi Gaia Scholars Research Fund), Kuaishou, and Bytedance.